1 1/2 + 5 y/3 = 7 1/6
Answer by stanbon(75887)

1 1/2 + 5 y/3 = 7 1/6
(3/2) + (5/3)y = (43/6)
----
Multiply thru by 6 to get;
9 + 10y = 43
10y = 32
y = 16/5
y = 3 1/5
================
Cheers,
Stan H.
================